OpenStudy (anonymous):

Find all the solutions to the equation Cos=-1/2 Please Help!

rvc (rvc):

cos(? )= 1/2

OpenStudy (anonymous):

huh?

OpenStudy (luigi0210):

You're missing your variable :)

OpenStudy (anonymous):

Im at a complete loss at how to go about this equation...

OpenStudy (luigi0210):

Are you sure it's \(\cos=-\frac{1}{2} \) and not \(\cos \theta=-\frac{1}{2} \) or something like that?

OpenStudy (anonymous):

Opps, I meant Cos x= -1/2 lol

OpenStudy (luigi0210):

Do you know how to start?

OpenStudy (anonymous):

No....